Circular No. 6185

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995S IN NGC 4456 R. H. McNaught, Anglo-Australian Observatory, reports his discovery

of a supernova (mag ∼ 17.5) on an R survey film taken by K. S. Russell with the U.K. Schmidt telescope (UKST) on July 3.4 UT. A confirming unfiltered CCD image was obtained by G. J. Garradd and McNaught with the 1.0-m reflector at Siding Spring on July 6. SN 1995S is located at α = 12^{h}27^{m}53^{s}.30, δ = −30^{o}05′55′′.4 (equinox 2000.0, uncertainty 0′′.4 in each coordinate), which is 12′′ east and 3′′ south of the center of NGC 4456. SN 1995S is offset from a star of mag 20 (clearly present on the ESO R and SERC J surveys) by ∼ 7′′ in p.a. 200^{o}; this star and another further south both lie on a spiral arm, and the supernova lies inside this arm. SN 1995S is not present on UKST R plates taken by Russell on Feb. 7 (limiting mag 21) and May 28 (limiting mag 20), nor on an I survey plate taken on Mar. 11 by M. J. Drinkwater; likewise, no image appears at the given position on several ESO and SERC B, J, R, and I survey films. A nearby star of mag ∼ 15.5 has position end figures 50^{s}.88, 06′08′′.0.

PKS 1622–297 J. D. Kurfess, J. E. Grove, K. McNaron-Brown, and W. N. Johnson,

Naval Research Laboratory, on behalf of the OSSE team; J. R. Mattox, University of Maryland and Universities Space Research Association; S. Wagner, Landessternwarte Heidelberg; and J. R. Webb, Florida International University, communicate: “Optical observations of PKS 1622–297 at Calar Alto, at La Silla, and with the SARA telescope on Kitt Peak indicate that PKS 1622–297 is now at mag ∼ 17 (at least 3 mag brighter than its historical value). The OSSE instrument aboard Compton detects this quasar from 50–200 keV at a significance of 6σ during the interval June 30–July 6, with an indication that the flux is somewhat higher on July 5– 6. The EGRET instrument aboard the Compton Observatory continues to detect PKS 1622–297. The high-energy γ-ray flux peaked on June 25 and has declined since then. The Compton observations of PKS 1622–297 were extended until July 10, and the OSSE instrument will continue monitoring the source through July 24. Observations at other wavelengths continue to be very important.”

1995 July 11 Daniel W. E. Green